OpenStudy (anonymous):

You have to prepare a pH 5.00 buffer, and you have the following 0.10M solutions available: HCOOH, HCOONa, CH3COOH, CH3COONa, HCN, and NaCN. A. The solutions you would see are CH3COOH, and CH3COONa... B. How many millileters of each solution would you use to make approximately a liter of the buffer? Can someone please help me with letter B? I don't know how to approach this problem.

OpenStudy (anonymous):

@CausticSyndicalist do you know?

OpenStudy (abb0t):

You can use \(\sf C_1V_1 = C_2V_2\) Where C = concentration and V = volume the 1 represents initial (what you're starting with) and 2 represents the final volume/concentration you want. Keep in mind that concentration is usually in units \(\sf \frac{mol}{L}\)

OpenStudy (anonymous):

There is no starting volume listed...how would we do that then

OpenStudy (anonymous):

@abb0t

OpenStudy (aaronq):

Start with the Henderson Hasselbalch equation: \(pH=pKa+log\dfrac{[A^-]}{[HA]}\)

OpenStudy (anonymous):

ok so far i have \[5.00=4.74 +\log(CH _{3}COO ^{-}/CH _{3}COOH)\]

OpenStudy (anonymous):

then you would subtract the 4.74 over and get \[.26=\log(CH _{3}COO ^{-}/CH _{3}COOH)\] and to get rid of the log you do \[10^{.26}\] which equals \[1.82\] then you will have \[1.82=(CH _{3}COO ^{-}/CH _{3}COOH)\] and this is where i'm stuck

OpenStudy (aaronq):

\(\dfrac{[A^-]}{[HA]}=\dfrac{1.82}{1} \) the total moles (assuming 1 L) are: \(M=\dfrac{n}{L}\rightarrow n=0.1~M*1L=0.1`moles\) so in terms of percent: \([A^-]=\dfrac{1.82}{1+1.82}*0.1~moles\) \([HA]=\dfrac{1}{1+1.82}*0.1~moles \) Find the volume representative of those moles

OpenStudy (anonymous):

A- would be .1 and HA would be .035 or 3.5*10^-2 ???

OpenStudy (aaronq):

moles of \(A^-\) are 0.64539 moles, therefore \(\sf 0.1~M=\dfrac{0.64539~moles}{L}\rightarrow L=0.6454~L ~or ~645.4 ~mL \)

OpenStudy (aaronq):

sorry it was 0.064539 moles

OpenStudy (anonymous):

ok but what would HA be then? sorry, i'm just pretty confused on this

OpenStudy (aaronq):

from above, the moles of HA are: \(\sf n_{HA}=\dfrac{1}{1+1.82}*0.1~moles=0.03546~moles\) then the volume is : \(\sf 0.1~M=\dfrac{0.03546~moles}{L}\rightarrow L=0.3546~L~or~345.6~mL\)